Each of two small spheres is charged positively, the total charge 52.6 micro C. Each sphere repelled from the other with force of 1.19 N. when the spheres are 1.94 m apart. Calculate the charge on each sphere.
1
Expert's answer
2020-02-06T09:25:07-0500

Let the charge of the first and second spheres "q_1" and "q_2". Their sum by the condition of the problem (1) "q_1+q_2=Q_{total}=52.6 \mu C"

The force of electrostatic repulsion is determined by the Coulomb's law as

(2) "F=k_e\frac{q_1\cdot q_2}{r^2}" , where "k_e=8.987\cdot 10^9 Nm^2C^{-2}" is Coulomb's constant [1].

Equations (1) and (2) are a system of two equations with two unknowns. Find solution of the system.

From (1)

(3) "q_2=Q_{total}-q_1"

Substitude (3) in (2) and making some simplifications and permutations we get the square eqution for determine "q_1"

(4) "q_1^2-Q_{total}\cdot q_1+\frac {F\cdot r^2}{k_e}=0"

The solution to this quadratic equation has the form

(5) "q_1=\frac{Q_{total}}{2}\pm \sqrt{(\frac{Q_{total}}{2})^2-\frac{F\cdot r^2}{k_e}}"

Determine all value of "q_1"

"\frac{Q_{total}}{2}=\frac{52.6\mu C}{2}=26.3\cdot 10^{-6}C=2.63\cdot 10^{-5}C"

"(\frac{Q_{total}}{2}) ^2=6.92\cdot 10^{-10}C^2"

"\frac{F\cdot r^2}{k_e}=\frac{1.19N\cdot (1.94m)^2}{8.987\cdot 10^9 Nm^2C^{-2}}=0.5\cdot 10^{-9}C^2=5\cdot 10^{-10}C^{2}"

"q_1=(2.63\cdot 10^{-5}\pm \sqrt{6.92-5}\cdot 10^{-5})C= \begin{cases}
 4.02 &\text{if }+ \\
 1.24 &\text{if } -
\end{cases}10^{-5}C=\begin{cases}
 40.2 &\text{if }+ \\
 12.4 &\text{if } -
\end{cases}\mu C"

Easy to see that

"q_2= \begin{cases}
 12.4 &\text{if }+ in\ eqution (5) \\
 40.2 &\text{if } - in\ equation (5)
\end{cases}\mu C"

So we actually have one solution namely, one of the spheres (anyway, which one) has a charge "40.2 \mu C" and the other "12.4 \mu C".

Answer: The charge on one sphere is 40.2 micro C and the other 12.4 micro C.

[1]https://en.wikipedia.org/wiki/Coulomb%27s_law